OpenStudy (anonymous):

Simplify (t2+6t)/(t-1)- (7)/(t-1)

OpenStudy (anonymous):

Is it the question?\[\frac{ t^{2}+6t }{ t-1 }-\frac{ 7 }{ t-1 }\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\frac{ t^2+6t-7 }{ t-1 }=\frac{ (t-1)(t+7) }{ t-1 }=t+7\]

OpenStudy (anonymous):

any question?

OpenStudy (anonymous):

why is (t-1) on the right side of the equation

OpenStudy (anonymous):

You can put it on the left since it product.

OpenStudy (anonymous):

so the final answer would be t+7 or t+7/t-1